sqrt(3x+10)=5-2x
solve for x
answer key says 3/4
my work
sqrt(3x+10)=5-2x
sqrt(3x+10)^2=(5-2x)^2
3x+10=25-4x
3x+4x+10=25-4x+4x
7x+10=25
7x+10-10=25-10
7x=15
7x/7=15/7
x=15/7

Answer by checkley77(12844)

sqrt(3x+10)=5-2x SQUARE BOTH SIDES.
[sqrt(3x+10)]^2=(5-2x)^2
3x+10=(25-20x+4X^2}
3x+10=25-20x+4X^2 COMBINE LIKE TERMS.
-4X^2+23x+10-25=0 MULTIPLY BY -1.
4X^2-23x+15=0
(4X-3)(X-5)=0
4X-3=0
4X=3
X=3/4 ANS.
X-5=0
X=5 ANS.